Add & Simplify
1/1+cos(x)+1/1-cos(x)
1
Expert's answer
2011-11-17T12:31:32-0500
1/1+cos(x)+1/1-cos(x)=(1+cos(x)+1-cos(x))/(1-(cos(x))^2)=2/(sin(x))^2=2(1+(cotan(x)^2))